# Asymptotic Velocity Domination

### v1

@@ 4.2 Hadamard and time consistent states
Both the Bunch–Davies vacuum (BD) and State of Low Energy (SLE) are also time consistent. Technically, their Bogoliubov parameters $(\lambda_{n},\mu_{n})$ are independent of the reference time $\tau_{0}$, albeit for distinct reasons.
@@
time-consistentHadamardBDSLE
Figure 4: Intersection of time-consistent and Hadamard states.
Using the notion of time consistent states as introduced in Subsection 3.2 the AVD result from subsection 3.4 can be paraphrased as follows: all time consistent states give rise to (matrix-) two-point functions that are asymptotically velocity dominated in the sense of (LABEL:avdresult). These states include many Hadamard states, in particular the Bunch-Davies vacuum and the States of Low Energy.
4.3 Quantum $\sigma$-field
@@
This is the remnant of the original constraints that still needs to be incorporated into the solution for $\phi$. In the quantum theory, one must define the field $\sigma$—that is, the components $T_{00}$ and $T_{01}$—as composite operators satisfying the partial conservation law and the constraint (4.20). An adiabatic renormalization of $T_{00}$ has been considered in [11].
@@
Hadamard states are tailored towards renormalization of quantum field theories in curved spacetimes [23, 24]. In the standard setting with a spatial slices diffeomorphic to $\mathbb{R}^{d}$ one can render the two-point function finite by subtracting its singular “Hadamard parametrix” in position space—i.e. the part singular on the lightcone—from the full two-point function. For spatially homogeneous backgrounds subtraction of a truncated Hadamard parametrix is consistent with the older adiabatic renormalization, see e.g. [26]. The precise relation has been elucidated in [22], see also Appendix A of [25].
In the polarized Gowdy model the spatial slices are circles, so the light-cone singularity acquires infinitely many $2\pi$–periodic images, see (4.5). Therefore, it is more convenient to work in Fourier space and parameterize the two-point function using its initial data $Z_{n}(\tau_{0})$. Since Hadamard states are singled out by the universal large $n$ asymptotics (4.4) of $|T_{n}(\tau)|^{2}$, it follows that the initial–value matrix $Z_{n}(\tau_{0})$ itself admits a universal large $n$ expansion for every Hadamard state. Explicitly,
@@ 5.  Conclusions
The immediate extension is of course to the non-polarized Gowdy system. A formal expansion of classical solutions around those of the velocity dominated system has been obtained in [29]. The construction of classical Dirac observables appears feasible by taking advantage of an underlying Lax pair. The required spatial periodicity (for the $T^{3}$ topology) however presents considerable complications compared to other two-Killing vector reductions. Through a Riemannian sigma-model formulation the quantum theory is amenable to an all-order perturbative (weak Newton coupling) analysis [8, 9, 10]. The quantum AVD property, on the other hand, is likely to be related to an expansion in inverse powers of Newton’s constant (strong coupling) and requires a different methodology.
Several issues in the polarized case also remained unexplored here. One is the fate of the curvature singularity as probed by the blow-up of the Kretschmann scalar, see (2.19). Among the reservations we have concerning this is that ${\cal R}_{2}/n$ is presumably not the integrand of a Dirac observable and that a quantum one-point function studied in [11, 17] is highly prescription dependent, while ${\cal R}_{2}$’s two-point function may not address the classical intuition. Similar remarks apply to the two-point function of the (exponentiated) $\sigma$-field. A functional integral quantization on the full (rather than reduced) phase space, including gauge fixing and Faddeev-Popov determinants, may be more fruitful with regard to extensions.
@@ Appendix A The Hamiltonian action and its VD counterpart
It is a general property of two-Killing vector reductions that entering with (2.1) into the Einstein-Hilbert action will produce a valid reduced action principle, i.e. one whose variations produce field equation coinciding with the ones obtained by directly specializing the Einstein field equations to metrics of the form (2.1). We therefore take the reduced Hamiltonian action of the polarized $T^{3}$ system as a starting point. The velocity dominated system is normally only defined through its field equations, see [3, 5]. In preparation for the quantum theory we develop here also an off-shell formulation in terms of an Hamiltonian action principle and its symmetries.
@@ A.1 Gowdy cosmologies: the Hamiltonian action and its symmetries
The field equations of the system are readily obtained by varying $S^{{\textsc{h}}}$. In addition to the constraints $\delta S^{{\textsc{h}}}/\delta n=-{\cal H}_{0}$, $\delta S^{{\textsc{h}}}/\delta s=-{\cal H}_{1}$, there are the velocity-momentum relations $e_{0}(\rho)+\lambda_{{\textsc{n}}}n\pi^{\sigma}=0$, $e_{0}(\sigma)+\lambda_{{\textsc{n}}}n\pi^{\rho}=0$, $\rho e_{0}(\phi)-\lambda_{{\textsc{n}}}n\pi^{\phi}=0$, and the evolution equations
@@ Appendix B Classical Dirac observables
Two Killing vector reductions of Einstein gravity typically admit an infinite set of nonlocal conserved charges whose existence is related to that of a Lax pair. On general grounds such conserved charges are expected to give rise to (normally elusive) Dirac observables, Poisson commuting with the constraints. For the polarized Gowdy system the conserved charges become linear functionals of the basic Gowdy scalar. A systematic characterization of Dirac observables for the polarized Gowdy system has been given by Torre [15], using an extended phase space formulation. Due to the use of a polar decomposition in the last step it falls however slightly short of providing fully explicit expressions for them. Here we provide such a construction on the original rather than an extended phase space in a way that makes contact to the velocity dominated limit and in principle generalizes to the non-polarized case.
A Dirac observable $\cal O$} is by definition a functional of the phase space variables that weakly Poisson commutes with the constraints
$\{{\cal H}_{0},{\cal O}\}\big{|}_{{\cal H}_{0}=0={\cal H}_{1}}=0=\{{\cal H}_{1},{\cal O}\}\big{|}_{{\cal H}_{0}=0={\cal H}_{1}}\,.$ (B.1)
@@
The goal in the following is to construct an infinite set of Dirac observables which satisfy (B.1), (B.2) strongly, off-shell, and without gauge fixing. Since ${\cal H}_{1}$ generates infinitesimal spatial translations spatial integrals of spatial +1+1 densities will satisfy the second condition in (B.1). A solution of the first condition, on the other hand, amounts to the construction of (nonlocal) conserved charges, which is normally elusive as it requires a constructive approach to the solution of the initial value problem. As mentioned, this is often feasible for the two Killing vector reductions of Einstein gravity. Moreover, for the polarized Gowdy system the construction simplifies and gives rise to an infinite set of conserved charges linear in the $\phi$. A streamlined approach is as follows.
@@
$\jmath_{0}^{\textsc{h}}:=\pi^{\phi}\,,\quad\jmath_{1}^{\textsc{h}}:=\frac{\rho n}{\lambda_{{\textsc{n}}}}\partial_{1}\phi\,.$ (B.3)
@@
Next, we aim at a one-parameter generalization of the form [^6] The origin of these expressions lies in the $\partial/\partial\theta$ derivatives of a Lax pair for the 2-Killing vector reductions. Use of the Lax pair is overkill for the polarized case but (B) turns out to generalize to the non-polarized Gowdy system.
$\jmath_{0}^{\textsc{h}}(\theta)$ $:=$ $\frac{1}{[\lambda_{{\textsc{n}}}^{2}(\theta+\tilde{\rho})^{2}-\rho^{2}]^{3/2}}[\lambda_{{\textsc{n}}}(\theta+\tilde{\rho})\pi^{\phi}-\frac{\rho^{2}}{\lambda_{{\textsc{n}}}}\partial_{1}\phi]\,$,
$\jmath_{1}^{\textsc{h}}(\theta)$ $:=$ $\frac{1}{[\lambda_{{\textsc{n}}}^{2}(\theta+\tilde{\rho})^{2}-\rho^{2}]^{3/2}}[(\theta+\tilde{\rho})\rho n\partial_{1}\phi-\rho n\pi^{\phi}]\,$, (B.5)
@@
and treat $\tilde{\rho}(x^{0},y^{1})$ as independent of $\rho$. For its Hamiltonian gauge variation one can self-consistently take
@@
Note that compared to $\delta_{\epsilon}^{\textsc{h}}\rho$ in (A.4) the roles of the descriptors $\epsilon,\epsilon^{1}$ are swapped. Using only (A.4) and (B.7) one finds
@@
Here we suppressed the shared time argument and assumed as before that the descriptors $\epsilon,\epsilon^{1}$ are spatially periodic. The basic idea now it to replace $\jmath_{0}^{\textsc{h}}(x,\theta)$ on the left hand side with its $2\pi$-periodic extension, i.e. with $\sum_{n}\jmath_{0}^{\textsc{h}}(x\!+\!2\pi n,\theta)$. Assuming that this sum converges, it will give rise on the right hand side to a telescopic sum which vanishes. In principle, the plain periodic extension can be modified by multiplying the $n$-th term by a function $s_{n}$ that is itself gauge invariant and obeys $s_{n}|_{x^{1}+2\pi}=s_{n+1}$ as a function of $x^{1}$. Since the only source of non-periodicity in our context is the combination $\theta+\tilde{\rho}$ it is natural to take $s_{n}$ a function of $\theta+\tilde{\rho}$, in which case $\delta^{\textsc{h}}_{\epsilon}s_{n}=0$ is a stringent requirement.
@@
$c_{0}(\tilde{\rho}+\theta,\rho)$ $:=$ $\sum_{n\in\mathbb{Z}}\frac{s_{n}(\theta\!+\!\tilde{\rho})\lambda_{{\textsc{n}}}(\theta+\tilde{\rho}-n\pi_{0}^{\sigma})}{[\lambda_{{\textsc{n}}}^{2}(\theta+\tilde{\rho}-n\pi_{0}^{\sigma})^{2}-\rho^{2}]^{3/2}}\,$,
$c_{1}(\tilde{\rho}+\theta,\rho)$ $:=$ $\sum_{n\in\mathbb{Z}}\frac{s_{n}(\theta\!+\!\tilde{\rho})}{[\lambda_{{\textsc{n}}}^{2}(\theta+\tilde{\rho}-n\pi_{0}^{\sigma})^{2}-\rho^{2}]^{3/2}}\,.$ (B.12)
For reasons that will become clear shortly we take for $s_{n}$ the generalized sign function $s_{n}(z-in\pi_{0}^{\sigma})=\exp\{3i{\rm Arg}(z-in\pi_{0}^{\sigma})\}(\exp\{-2i{\rm Arg}(z-in\pi_{0}^{\sigma})\})^{3/2}$, ${\rm Im}z\neq 0$. The sums (B) then converge absolutely and for ${\rm Im}\theta\neq 0$ define spatially periodic functions of $x^{0},x^{1}$. The summed expansions (LABEL:csums2) below also show them to be smooth in $\rho,\tilde{\rho}$ for ${\rm Im}(\theta)\neq 0$. Inserting (B) into (B) gives
$\jmath_{0}^{\textsc{p}}(\theta):=\sum_{n\in\mathbb{Z}}s_{n}(\theta+\tilde{\rho})\jmath^{\textsc{h}}_{0}(\theta-n\pi_{0}^{\sigma})=c_{0}(\tilde{\rho}+\theta,\rho)\pi^{\phi}-c_{1}(\tilde{\rho}+\theta,\rho)\frac{\rho^{2}}{\lambda_{{\textsc{n}}}}\partial_{1}\phi\,$,
$\jmath_{1}^{{\textsc{p}}}(\theta):=\sum_{n\in\mathbb{Z}}s_{n}(\theta+\tilde{\rho})\jmath^{\textsc{h}}_{1}(\theta-n\pi_{0}^{\sigma})=c_{0}(\tilde{\rho}+\theta,\rho)\frac{\rho n}{\lambda_{{\textsc{n}}}}\partial_{1}\phi-c_{1}(\tilde{\rho}+\theta,\rho)\rho n\pi^{\phi}\,.$ (B.13)
The condition for the summed versions to obey the conservation equation (B.8) amounts to a simple pair of differential equations for $c_{0},c_{1}$,
$\partial_{\tilde{\rho}}(\rho^{2}c_{1})=-\lambda_{{\textsc{n}}}\rho\partial_{\rho}c_{0}\,,\quad\rho\partial_{\tilde{\rho}}c_{0}=-\lambda_{{\textsc{n}}}\partial_{\rho}(\rho^{2}c_{1})\,.$ (B.14)
@@
$\!\!\!\!\!\!\!\!\!\!\frac{\lambda_{{\textsc{n}}}(\theta\!+\!\tilde{\rho}\!-\!n\pi_{0}^{\sigma})}{[\lambda_{{\textsc{n}}}^{2}(\theta\!+\!\tilde{\rho}\!-\!n\pi_{0}^{\sigma})^{2}-\rho^{2}]^{3/2}}=\sum_{l\geq 0}{-3/2\choose l}(-\rho^{2})^{l}s_{n}(\theta\!+\!\tilde{\rho})[\lambda_{{\textsc{n}}}(\theta\!+\!\tilde{\rho}\!-\!n\pi_{0}^{\sigma})]^{-2l-2}$
$\!\!\!\!\!\!\!\!\!\!\frac{1}{[\lambda_{{\textsc{n}}}^{2}(\theta\!+\!\tilde{\rho}\!-\!n\pi_{0}^{\sigma})^{2}-\rho^{2}]^{3/2}}=\sum_{l\geq 0}{-3/2\choose l}(-\rho^{2})^{l}s_{n}(\theta\!+\!\tilde{\rho})[\lambda_{{\textsc{n}}}(\theta\!+\!\tilde{\rho}\!-\!n\pi_{0}^{\sigma})]^{-2l-3}\,$, (B.15)
@@
$\!\!\!\!\!\!\!\!\!\!c_{0}(\theta\!+\!\tilde{\rho},\rho)$ $\!\!=\!\!$ $\Big{(}\frac{\pi}{\lambda_{{\textsc{n}}}\pi_{0}^{\sigma}}\Big{)}^{2}\sum_{l\geq 0}{-3/2\choose l}\frac{(-\rho^{2}/\lambda_{{\textsc{n}}}^{2})^{l}}{(2l+1)!}\Big{(}\frac{\partial}{\partial\theta}\Big{)}^{2l}\!\frac{1}{\sin^{2}\big{(}\frac{\pi}{\pi_{0}^{\sigma}}(\theta\!+\!\tilde{\rho})\big{)}}$,
$\!\!\!\!\!\!\!\!\!\!c_{1}(\theta\!+\!\tilde{\rho},\rho)$ $\!\!=\!\!$ $-\frac{1}{\lambda_{{\textsc{n}}}}\Big{(}\frac{\pi}{\lambda_{{\textsc{n}}}\pi_{0}^{\sigma}}\Big{)}^{2}\sum_{l\geq 0}{-3/2\choose l}\frac{(-\rho^{2}/\lambda_{{\textsc{n}}}^{2})^{l}}{(2l+2)!}\Big{(}\frac{\partial}{\partial\theta}\Big{)}^{2l+1}\!\frac{1}{\sin^{2}\big{(}\frac{\pi}{\pi_{0}^{\sigma}}(\theta\!+\!\tilde{\rho})\big{)}}$. (B.17)
As a check on the innocuous nature of the $s_{n}$ modulation in (B) one can verify that the expressions (B) indeed solve the differential equations (B.14). Hence
${\cal O}(\theta):=\int_{0}^{2\pi}\!\frac{dx^{1}}{2\pi}\big{\{}c_{0}(\tilde{\rho}+\theta,\rho)\pi^{\phi}-c_{1}(\tilde{\rho}+\theta,\rho)\frac{\rho^{2}}{\lambda_{{\textsc{n}}}}\partial_{1}\phi\big{\}}\,,\quad{\rm Im}\theta\neq 0\,$, (B.18)
is a one-parameter family of off-shell Dirac observables for the polarized $T^{3}$ Gowdy cosmologies. Reading $\rho$ as ‘time’ the series (B) for $c_{0},c_{1}$ is especially natural in the context of AVD, see below.
For later use we note a third rewriting of the $c_{0},c_{1}$. Using
@@
$c_{0}(\theta+\tilde{\rho},\rho)$ $\!\!=\!\!$ $-\Big{(}\frac{2\pi}{\lambda_{{\textsc{n}}}\pi_{0}^{\sigma}}\Big{)}^{2}\sum_{n\geq 1}ne^{\pm 2i\pi n(\theta+\tilde{\rho})/\pi_{0}^{\sigma}}J_{0}\Big{(}\frac{2n\rho\pi}{\lambda_{{\textsc{n}}}\pi_{0}^{\sigma}}\Big{)}\,$,
$c_{1}(\theta+\tilde{\rho},\rho)$ $\!\!=\!\!$ $\pm i\Big{(}\frac{2\pi}{\pi_{0}^{\sigma}\lambda_{{\textsc{n}}}}\Big{)}^{2}\sum_{n\geq 1}ne^{\pm 2i\pi n(\theta+\tilde{\rho})/\pi_{0}^{\sigma}}\;\frac{1}{\rho}J_{1}\Big{(}\frac{2n\rho\pi}{\lambda_{{\textsc{n}}}\pi_{0}^{\sigma}}\Big{)}\,$, (B.20)

### v2

@@ 4.2 Hadamard and time consistent states
Both the Bunch–Davies vacuum (BD) and State of Low Energy (SLE) are also time consistent. Technically, this is because their Bogoliubov parameters $(\lambda_{n},\mu_{n})$ are independent of the reference time $\tau_{0}$, albeit for distinct reasons.
@@
time‐consistentHadamardBDSLE
Figure 4: Intersection of time‐consistent and Hadamard states.
Using the notion of time consistent states as introduced in Subsection 3.2 the AVD Results 3.2, 3.3 can be paraphrased as follows: all time consistent states give rise to (matrix-) two-point functions that are asymptotically velocity dominated in the sense of (LABEL:avdresult). These states include many Hadamard states, in particular the Bunch-Davies vacuum and the States of Low Energy.
4.3 Quantum $\sigma$‑field
@@
This is the remnant of the original constraints that still needs to be incorporated into the solution for $\phi$. In the quantum theory, one must define the field $\sigma$—that is, the components $T_{00}$ and $T_{01}$—as composite operators satisfying the partial conservation law and the constraint (4.20). An adiabatic renormalization [28] of $T_{00}$ has been considered early on in [11].
@@
Hadamard states are tailored towards renormalization of quantum field theories in curved spacetimes [25, 26]. In the standard setting with a spatial slices diffeomorphic to $\mathbb{R}^{d}$ one can render the two‐point function finite by subtracting its singular “Hadamard parametrix” in position space—i.e. the part singular on the lightcone—from the full two‐point function. For spatially homogeneous backgrounds subtraction of a truncated Hadamard parametrix is consistent with the older adiabatic renormalization, see e.g. [29]. The precise relation has been elucidated in [24], see also Appendix A of [27].
In the polarized Gowdy model the spatial slices are circles, so the light‐cone singularity acquires infinitely many $2\pi$–periodic images, see (4.5). Therefore, it is more convenient to work in Fourier space and parameterize the two‐point function using its initial data $Z_{n}(\tau_{0})$. Since Hadamard states are singled out by the universal large $n$ asymptotics (4.4) of $|T_{n}(\tau)|^{2}$, it follows that the initial–value matrix $Z_{n}(\tau_{0})$ itself admits a universal large $n$ expansion for every Hadamard state. Explicitly,
@@ 5.  Conclusions
The immediate extension is of course to the non-polarized Gowdy system. A formal expansion of classical solutions around those of the velocity dominated system has been obtained in [32]. The construction of classical Dirac observables is feasible by taking advantage of an underlying Lax pair [33]. The required spatial periodicity (for the $T^{3}$ topology) however presents considerable complications compared to other two-Killing vector reductions. Through a Riemannian sigma-model formulation the quantum theory is amenable to an all-order perturbative (weak Newton coupling) analysis [8, 9, 10]. The quantum AVD property, on the other hand, is likely to be related to an expansion in inverse powers of Newton’s constant (strong coupling) and requires a different methodology.
@@ Appendix A The Hamiltonian action and its VD counterpart
It is a general property of two-Killing vector reductions that entering with (2.1) into the Einstein-Hilbert action will produce a valid reduced action principle, i.e. one whose variations produce field equation coinciding with the ones obtained by directly specializing the Einstein field equations to metrics of the form (2.1). We therefore take the reduced Hamiltonian action of the polarized $T^{3}$ system as a starting point, see [13, 15]. The velocity dominated system is normally only defined through its field equations, see [3, 5]. In preparation for the quantum theory we develop here also an off-shell formulation in terms of an Hamiltonian action principle and its symmetries.
@@ A.1 Gowdy cosmologies: the Hamiltonian action and its symmetries
Unless stated otherwise we take the descriptors $\epsilon,\epsilon^{1}$ to be spatially periodic. The field equations of the system are readily obtained by varying $S^{{\textsc{h}}}$. In addition to the constraints $\delta S^{{\textsc{h}}}/\delta n=-{\cal H}_{0}$, $\delta S^{{\textsc{h}}}/\delta s=-{\cal H}_{1}$, there are the velocity-momentum relations $e_{0}(\rho)+\lambda_{{\textsc{n}}}n\pi^{\sigma}=0$, $e_{0}(\sigma)+\lambda_{{\textsc{n}}}n\pi^{\rho}=0$, $\rho e_{0}(\phi)-\lambda_{{\textsc{n}}}n\pi^{\phi}=0$, and the evolution equations
@@ Appendix B Classical Dirac observables
Two Killing vector reductions of Einstein gravity often admit an infinite set of nonlocal conserved charges whose existence is related to that of a Lax pair. On general grounds such conserved charges are expected to give rise to (normally elusive) Dirac observables, Poisson commuting with the constraints. For the polarized Gowdy system the conserved charges become linear functionals of the basic Gowdy scalar. A systematic characterization of Dirac observables for the polarized Gowdy system has been given by Torre [15], using an extended phase space formulation. Due to the use of a polar decomposition in the last step it falls however slightly short of providing fully explicit expressions for them. Here we provide such a construction on the original rather than an extended phase space in a way that makes contact to the velocity dominated limit and in principle generalizes to the non-polarized case.
A Dirac observable $\cal O$} is by definition a functional of the phase space variables that weakly Poisson commutes with the constraints, see e.g. [16]. In the situation at hand, we aim at the strong vanishing
$\{{\cal H}_{0},{\cal O}\}=0=\{{\cal H}_{1},{\cal O}\}\,.$ (B.1)
@@
The goal in the following is to construct an infinite set of Dirac observables which satisfy (B.1), (B.2) strongly, off-shell, and without gauge fixing. See e.g. [16] for the significance of the latter. Since ${\cal H}_{1}$ generates infinitesimal spatial translations spatial integrals of spatial +1+1 densities will satisfy the second condition in (B.1). A solution of the first condition, on the other hand, necessitates to the construction of (nonlocal) conserved charges, which is normally elusive as it requires a constructive approach to the solution of the initial value problem. As mentioned, this is often feasible for the two Killing vector reductions of Einstein gravity, but for Gowdy cosmologies has not been accomplished beyond the polarized sector. In the following we describe a construction of off-shell Dirac observables for the polarized Gowdy system that generalizes to the unpolarized sector.
@@
$\jmath_{0}^{\textsc{h}}:=\pi^{\phi}\,,\quad\jmath_{1}^{\textsc{h}}:=\frac{\rho n}{\lambda_{{\textsc{n}}}}\partial_{1}\phi\,$, (B.3)
@@
Next, we aim at a one-parameter generalization of the form [^6] The origin of these expressions lies in the $\partial/\partial\theta$ derivatives of a Lax pair for the 2-Killing vector reductions, see [33]. Use of the Lax pair is overkill for the polarized case but (B) turns out to generalize to the non-polarized Gowdy system.
$\jmath_{0}^{\textsc{h}}(\theta)$ $:=$ $\frac{\lambda_{{\textsc{n}}}^{2}}{[\lambda_{{\textsc{n}}}^{2}(\theta+\tilde{\rho})^{2}-\rho^{2}]^{3/2}}[\lambda_{{\textsc{n}}}(\theta+\tilde{\rho})\pi^{\phi}-\frac{\rho^{2}}{\lambda_{{\textsc{n}}}}\partial_{1}\phi]\,$,
$\jmath_{1}^{\textsc{h}}(\theta)$ $:=$ $\frac{\lambda_{{\textsc{n}}}^{2}}{[\lambda_{{\textsc{n}}}^{2}(\theta+\tilde{\rho})^{2}-\rho^{2}]^{3/2}}[(\theta+\tilde{\rho})\rho n\partial_{1}\phi-\rho n\pi^{\phi}]\,$, (B.5)
@@
for some reference point $y^{1}$. For its Hamiltonian gauge variation one can self-consistently take
@@
This comes about as follows: since $\tilde{\rho}$ is not spatially periodic (see below) the natural class of gauge-type spatial diffeomorphisms for its are those of the real line vanishing outside a compact region. Evaluating the gauge variation of the integral in (B.7) one finds that that contribution from the $y^{1}$ endpoint vanishes if $y^{1}\rightarrow$-$\infty$. One can therefore consistently take $\tilde{\rho}(x^{0},-\infty)=0$ and obtains (B.7). Once (B.7) is in place we treat $\tilde{\rho}$ as spatially quasi-periodic via (B.9) and for the fundamental domain $0,2\pi]$ use gauge descriptors of the same type as in (A.4).
Using only (A.4) and (B.7) one finds
@@
Here we suppressed the shared time argument and assumed as before that the descriptors $\epsilon,\epsilon^{1}$ are spatially periodic. The basic idea now is to replace $\jmath_{0}^{\textsc{h}}(x,\theta)$ on the left hand side with its $2\pi$-periodic extension, i.e. with $\sum_{n}\jmath_{0}^{\textsc{h}}(x\!+\!2\pi n,\theta)$. Assuming that this sum converges, it will give rise on the right hand side to a telescopic sum which vanishes. In principle, the plain periodic extension can be modified by multiplying the $n$-th term by a function $s_{n}$ that is itself gauge invariant and obeys $s_{n}|_{x^{1}+2\pi}=s_{n+1}$ as a function of $x^{1}$. Since the only source of non-periodicity in our context is the combination $\theta+\tilde{\rho}$ it is natural to take $s_{n}$ a function of $\theta+\tilde{\rho}$, in which case $\delta^{\textsc{h}}_{\epsilon}s_{n}=0$ is a stringent requirement.
@@
$c_{0}(\tilde{\rho}+\theta,\rho)$ $:=$ $\sum_{n\in\mathbb{Z}}\frac{s_{n}(\theta\!+\!\tilde{\rho})\lambda_{{\textsc{n}}}^{3}(\theta+\tilde{\rho}-n\pi_{0}^{\sigma})}{[\lambda_{{\textsc{n}}}^{2}(\theta+\tilde{\rho}-n\pi_{0}^{\sigma})^{2}-\rho^{2}]^{3/2}}\,$,
$c_{1}(\tilde{\rho}+\theta,\rho)$ $:=$ $\sum_{n\in\mathbb{Z}}\frac{\lambda_{{\textsc{n}}}^{3}s_{n}(\theta\!+\!\tilde{\rho})}{[\lambda_{{\textsc{n}}}^{2}(\theta+\tilde{\rho}-n\pi_{0}^{\sigma})^{2}-\rho^{2}]^{3/2}}\,.$ (B.12)
For reasons that will become clear shortly we take for $s_{n}$ the generalized sign function $s_{n}(z-in\pi_{0}^{\sigma})=\exp\{3i{\rm Arg}(z-in\pi_{0}^{\sigma})\}(\exp\{2i{\rm Arg}(z-in\pi_{0}^{\sigma})\})^{-3/2}$, ${\rm Im}z\neq 0$. The sums (B) then converge absolutely and for ${\rm Im}\theta\neq 0$ define spatially periodic functions of $x^{0},x^{1}$. The summed expansions (LABEL:csums2) below also show them to be smooth in $\rho,\tilde{\rho}$ for ${\rm Im}(\theta)\neq 0$. Inserting (B) into (B) gives
$\jmath_{0}^{\textsc{p}}(\theta):=\sum_{n\in\mathbb{Z}}s_{n}(\theta+\tilde{\rho})\jmath^{\textsc{h}}_{0}(\theta-n\pi_{0}^{\sigma})=c_{0}(\tilde{\rho}+\theta,\rho)\pi^{\phi}-c_{1}(\tilde{\rho}+\theta,\rho)\frac{\rho^{2}}{\lambda_{{\textsc{n}}}^{2}}\partial_{1}\phi\,$,
$\jmath_{1}^{{\textsc{p}}}(\theta):=\sum_{n\in\mathbb{Z}}s_{n}(\theta+\tilde{\rho})\jmath^{\textsc{h}}_{1}(\theta-n\pi_{0}^{\sigma})=c_{0}(\tilde{\rho}+\theta,\rho)\frac{\rho}{\lambda_{{\textsc{n}}}}n\partial_{1}\phi-c_{1}(\tilde{\rho}+\theta,\rho)\frac{\rho}{\lambda_{{\textsc{n}}}}n\pi^{\phi}\,.$ (B.13)
The condition for the summed versions to obey equation (B.8) amounts to a simple pair of differential equations for $c_{0},c_{1}$,
$\partial_{\tilde{\rho}}(\rho^{2}c_{1})=-\lambda_{{\textsc{n}}}^{2}\rho\partial_{\rho}c_{0}\,,\quad\rho\partial_{\tilde{\rho}}c_{0}=-\partial_{\rho}(\rho^{2}c_{1})\,.$ (B.14)
@@
$\!\!\!\!\!\!\!\!\!\!\frac{\lambda_{{\textsc{n}}}^{3}(\theta\!+\!\tilde{\rho}\!-\!n\pi_{0}^{\sigma})}{[\lambda_{{\textsc{n}}}^{2}(\theta\!+\!\tilde{\rho}\!-\!n\pi_{0}^{\sigma})^{2}-\rho^{2}]^{3/2}}=\sum_{l\geq 0}{-3/2\choose l}(-\rho^{2})^{l}s_{n}(\theta\!+\!\tilde{\rho})\lambda_{{\textsc{n}}}^{2}[\lambda_{{\textsc{n}}}(\theta\!+\!\tilde{\rho}\!-\!n\pi_{0}^{\sigma})]^{-2l-2},$
$\!\!\!\!\!\!\!\!\!\!\frac{\lambda_{{\textsc{n}}}^{3}}{[\lambda_{{\textsc{n}}}^{2}(\theta\!+\!\tilde{\rho}\!-\!n\pi_{0}^{\sigma})^{2}-\rho^{2}]^{3/2}}=\sum_{l\geq 0}{-3/2\choose l}(-\rho^{2})^{l}s_{n}(\theta\!+\!\tilde{\rho})\lambda_{{\textsc{n}}}^{3}[\lambda_{{\textsc{n}}}(\theta\!+\!\tilde{\rho}\!-\!n\pi_{0}^{\sigma})]^{-2l-3},$ (B.15)
@@
$\!\!\!\!\!\!\!\!\!\!c_{0}(\theta\!+\!\tilde{\rho},\rho)$ $\!\!=\!\!$ $\Big(\frac{\pi}{\pi_{0}^{\sigma}}\Big)^{2}\sum_{l\geq 0}{-3/2\choose l}\frac{(-\rho^{2}/\lambda_{{\textsc{n}}}^{2})^{l}}{(2l+1)!}\Big(\frac{\partial}{\partial\theta}\Big)^{2l}\!\frac{1}{\sin^{2}\big(\frac{\pi}{\pi_{0}^{\sigma}}(\theta\!+\!\tilde{\rho})\big)}$,
$\!\!\!\!\!\!\!\!\!\!c_{1}(\theta\!+\!\tilde{\rho},\rho)$ $\!\!=\!\!$ $-\Big(\frac{\pi}{\pi_{0}^{\sigma}}\Big)^{2}\sum_{l\geq 0}{-3/2\choose l}\frac{(-\rho^{2}/\lambda_{{\textsc{n}}}^{2})^{l}}{(2l+2)!}\Big(\frac{\partial}{\partial\theta}\Big)^{2l+1}\!\frac{1}{\sin^{2}\big(\frac{\pi}{\pi_{0}^{\sigma}}(\theta\!+\!\tilde{\rho})\big)}$. (B.17)
As a check on the innocuous nature of the $s_{n}$ modulation in (B) one can verify that the expressions (B) indeed solve the differential equations (B.14). This yields an expansion of $\jmath_{0}^{{\textsc{p}}}(\theta)$ of the form $\jmath_{0}^{{\textsc{p}}}(\theta)=\sum_{j\geq 0}\jmath_{0}^{{\textsc{p}}}(\theta)_{2j}$, where the $j$-th term has scaling weight $\ell^{-2j}$ under (A.11). Explicitly,
$\jmath_{0}^{{\textsc{p}}}(\theta)_{0}$ $:=$ $\Big(\frac{\pi}{\pi_{0}^{\sigma}}\Big)^{2}\frac{\pi^{\phi}}{\sin^{2}\big(\frac{\pi}{\pi_{0}^{\sigma}}(\theta+\tilde{\rho})\big)}\,$, (B.18)
$\jmath_{0}^{{\textsc{p}}}(\theta)_{2j}$ $:=$ $\Big(\frac{\pi}{\pi_{0}^{\sigma}}\Big)^{2}{-3/2\choose j}\frac{(-1)^{j}}{(2j\!+\!1)!}\Big(\frac{\rho^{2}}{\lambda_{{\textsc{n}}}^{2}}\Big)^{j}$
$\times\Big(\frac{\partial}{\partial\theta}\Big)^{2j}\Big[\frac{\pi^{\phi}}{\sin^{2}\big(\frac{\pi}{\pi_{0}^{\sigma}}(\theta+\tilde{\rho})\big)}+\frac{1}{(2j\!+\!2)}\frac{\partial}{\partial\theta}\frac{\partial_{1}\phi}{\sin^{2}\big(\frac{\pi}{\pi_{0}^{\sigma}}(\theta+\tilde{\rho})\big)}\Big$]\,.
Finally, we note a third rewriting of the $c_{0},c_{1}$. Using
@@
$c_{0}(\theta+\tilde{\rho},\rho)$ $\!\!=\!\!$ $-\Big(\frac{2\pi}{\pi_{0}^{\sigma}}\Big)^{2}\sum_{n\geq 1}ne^{\pm 2i\pi n(\theta+\tilde{\rho})/\pi_{0}^{\sigma}}J_{0}\Big(\frac{2n\rho\pi}{\lambda_{{\textsc{n}}}\pi_{0}^{\sigma}}\Big)\,$,
$c_{1}(\theta+\tilde{\rho},\rho)$ $\!\!=\!\!$ $\pm i\Big(\frac{2\pi}{\pi_{0}^{\sigma}}\Big)^{2}\sum_{n\geq 1}ne^{\pm 2i\pi n(\theta+\tilde{\rho})/\pi_{0}^{\sigma}}\;\frac{\lambda_{{\textsc{n}}}}{\rho}J_{1}\Big(\frac{2n\rho\pi}{\lambda_{{\textsc{n}}}\pi_{0}^{\sigma}}\Big)\,$, (B.20)